FTS (Gavin and Paolo) Matching the configuration today. The configuration of the pilot node has to be adapted. a modification for new updated needed. Next update of the pilot service forseen, depending on how far the fixes with the configuration are but probably next week (Gavin and Paolo). The pilot server will be update to the V1.4.1 while the servive node to the latest 1.3

oplaproc79 has been fixed (Vlado)

Fixes for the libs are over, it is pending therefore on the deployment. The RPMs will be joined to the repository. fir LCHb a prestaging has been done as temporary solution. CMS is working with a SE in Taiwan working with srmcopy. Commands have to be optimized. It has been communicated to the site manager and sent emails to CNAF and PIC to follow the same procedure (Olof)

Actions in all services to include the new kernel is foreseen

LHCb presented beautiful plots during the night transfers with a rate average of 50MB/s. (Roberto) more than 1000 files from tape has been transfered although this number is shared with PHEDEX transfers since they are together in production

-- Main.pmendez - 13 Oct 2005

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2005-10-13 - PatriciaMendezLorenzo
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback